### v1

# Forgetting is Everywhere

Abstract
A fundamental challenge in developing general learning algorithms is their tendency to forget past knowledge when adapting to new data. Addressing this problem requires a principled understanding of forgetting; yet, despite decades of study, no unified definition has emerged that provides insights into the underlying dynamics of learning. We propose an algorithm- and task-agnostic theory that characterises forgetting as a lack of self-consistency in a learner’s predictive distribution over future experiences, manifesting as a loss of predictive information. Our theory naturally yields a general measure of an algorithm’s propensity to forget. To validate the theory, we design a comprehensive set of experiments that span classification, regression, generative modelling, and reinforcement learning. We empirically demonstrate how forgetting is present across all learning settings and plays a significant role in determining learning efficiency. Together, these results establish a principled understanding of forgetting and lay the foundation for analysing and improving the information retention capabilities of general learning algorithms. Project page: https://ben-sanati.github.io/forgetting-is-everywhere-project/.
1 Introduction
Forgetting is a ubiquitous yet poorly understood phenomenon in machine learning (catainf). When a learner updates its beliefs based on new observations, it often forgets prior knowledge, leading to a degradation in performance on previous observations. Although this behaviour is well documented in continual learning (CL; overcomeCF, overcomeCF) and reinforcement learning (RL; atkinson2021pseudo, atkinson2021pseudo; khetarpal2022towards, khetarpal2022towards), it is not confined to non-stationary settings. Even under standard training regimes with independent and identically distributed (i.i.d.) data, models can forget previously acquired abilities, reducing learning efficiency and stability (chunking).
Classical learning theory offers limited tools to analyse forgetting. Concepts such as bias, variance, and generalisation describe performance at fixed times, typically used at convergence. Forgetting, by contrast, is inherently temporal; it concerns what a model retains, what it discards, and how it balances adaptation with stability over time (jagielski2022measuring).
Most existing studies of forgetting come from the CL literature, where metrics typically track performance degradation on earlier tasks after training on later tasks (chaudhry2018riemannian). Although these measures are widely adopted, they poorly capture forgetting and frequently conflate two distinct phenomena: backward transfer, where new learning improves performance on past tasks (benavides2022theory), and forgetting, where updates erode prior knowledge (jagielski2022measuring), making it challenging to distinguish between constructive and destructive adaptation.
In this paper, we introduce a new conceptual foundation for understanding forgetting, built on the following insight:
If a learner updates its predictions on data it already expects, such an update cannot represent the acquisition of new information; it must instead represent the loss of previously acquired knowledge.
This allows us to characterise forgetting in a precise and general manner. We ask: What is forgetting? When and why does it occur? How does forgetting impact learning?
We address these questions through the following contributions:
  1. 1.
    Formulation of learning over time: Inspired by hutter2005 and dong2022simple, we define a general theoretical formulation for reasoning about how learners acquire, retain, and lose capabilities during learning (§ 3).
  2. 2.
    Conceptualisation of forgetting: Building on our insight, we formulate forgetting from a predictive perspective: as a property of how a learner’s predictions evolve over time. Specifically, we define forgetting as a violation of self-consistency in the learner’s predictive distribution, providing a unified probabilistic foundation that generalises prior conceptions of forgetting (§ 4.2).
  3. 3.
    Operational measure: This conceptualisation naturally yields a measure of the propensity to forget, which we propose to assess the validity of our formalism (Definition 4.2).
  4. 4.
    Empirical characterisation: We empirically investigate the propensity to forget across diverse environments and learning paradigms, including regression, classification, generative modelling, CL, and RL. The results confirm that the observed forgetting dynamics conform to the expected theoretical characteristics of forgetting and further reveal a trade-off between forgetting and training efficiency, with deep RL being particularly susceptible to forgetting (§ 5).
2 Related work
In this paper, we study the prevalence of forgetting in general learning algorithms. While terminology varies across fields, all phenomena involving the loss of previously acquired knowledge reflect the same underlying process. For clarity, we refer to all such phenomena collectively as forgetting.
Forgetting in supervised settings.
Forgetting is often studied in the context of CL (de2021continual; wang2024comprehensive). In this literature, various proxy metrics are used to measure forgetting in settings where a learner encounters data sequentially from different distributions. Here, forgetting is calculated as the change in performance on previously observed distributions (overcomeCF; chaudhry2018efficient; chaudhry2018riemannian).
However, this approach cannot distinguish between two distinct effects: backwards transfer, where learning on new data improves performance on past tasks (benavides2022theory), and forgetting, where learning on new data degrades previously acquired knowledge (catainf; jagielski2022measuring). Performance on previous tasks depends on both how current data informs the learner about past data (backwards transfer) and the extent to which updating on new tasks results in a loss of previous knowledge. Furthermore, most metrics are limited to CL settings and do not generalise to all learners or domains, motivating the need for a more general definition of forgetting.
Various approaches have sought to quantify forgetting. In toy settings with known data distributions, forgetting can be measured relative to an oracle (lee2021continual). Others examine changes in internal representations (kim2025understanding) or model the trade-off between generalisation and forgetting (raghavan2021formalizing). Meanwhile, mechanistic accounts in associative memory models demonstrate how interference between correlated patterns leads to retrieval errors (hopfield_discrete; Amit). A related line of work studies models trained on their own generated outputs, typically in the context of generative model collapse or stability (alemohammad2023self; shumailov2023curse; bertrand2023stability; scholten2025model). These studies implicitly touch on the phenomenon we formalise as forgetting: updating a model on self-generated targets can degrade previously encoded knowledge or induce systematic shifts in its predictions.
Forgetting in RL.
Forgetting has also been studied in continual RL; however, it has received less systematic study than in supervised CL. Early discussions of continual RL already noted the risk of knowledge degradation when agents adapt over long time horizons (ring1994continual; ring1997child). More recent surveys highlight that catastrophic forgetting is a persistent challenge in RL, where continual adaptation to changing state distributions or reward signals can interfere with previously acquired knowledge (khetarpal2022towards; kumar2025continual).
Empirical studies document several forms of forgetting. Value-based methods with function approximation often lose performance on earlier estimates (mnih2015human; van2018deep). This is closely related to the phenomenon of policy churn identified by (schaul2022phenomenon) in which the greedy policy of a value-based learner changes in a large portion of the input space after just a few updates. Policy gradient methods are also prone to overwriting earlier strategies during continual adaptation (ring1994continual; kirkpatrick2017overcoming). Replay buffers are widely used to mitigate these effects by reintroducing past experiences during training (random; fedus2020revisiting). However, such methods only partially alleviate forgetting without addressing its root cause, and thus are insufficient solutions in settings like constrained continual RL (dong2022simple; kumar2025continual).
Unlike in supervised CL, where task boundaries are externally defined and forgetting is measured via performance drops on prior tasks, in RL, the data distribution evolves continuously with the learner’s own policy. As a result, forgetting may manifest as interference with previously learned predictions and the gradual disappearance of coverage over regions of the state space that the agent no longer visits (sutton1998reinforcement). Despite these differences, metrics for measuring forgetting in RL remain underdeveloped, and most approaches borrow from supervised CL.
In this paper, we propose a conceptualisation of forgetting that is broadly applicable across learning paradigms and provides a principled distinction between forgetting and related phenomena, such as backwards transfer. Making this distinction is crucial for understanding whether learner updates reflect knowledge loss or beneficial reuse of information. We further demonstrate the generality of our conceptualisation and show how it supports an empirical measure of forgetting that aligns with theoretical expectations.
3 Learning and inference processes
We present a general framework in which supervised learning, RL, and generative modelling are all specific cases of a single stochastic interaction process. The formalism draws inspiration from the agent-environment perspective from general RL (hutter2005; Lattimore2014; dong2022simple; Abel2023; hutter2024; kumar2025continual), but is adapted to emphasise conventions standard in machine learning (bishop2006pattern; goodfellow2016deep).
Notation.
We let capital calligraphic letters denote measurable spaces $(\mathcal{X})$ [^1] We assume that all measurable spaces $(\mathcal{X},\mathcal{Y},\mathcal{Z},\dots)$ are standard Borel (Borel $\sigma$-algebras of Polish spaces), such that conditional probability kernels $p(\cdot\mid h,y)$ and $q(\cdot\mid z,x)$ exist and are measurable., lowercase letters denote elements $(x)$ or functions $(f)$, and uppercase italics denote random variables $(X)$. For any measurable space $\mathcal{X}$, we let $\mathcal{D}(\mathcal{X})$ denote the set of probability distributions over $\mathcal{X}$. For example, a mapping
$p:\mathcal{X}\times\mathcal{Y}\to\mathcal{D}(\mathcal{Z})$ (1)
is interpreted as a family of conditional distributions $p(\cdot\mid x,y)$ on $\mathcal{Z}$, indexed by $(x,y)\in\mathcal{X}\times\mathcal{Y}$.
3.1 Preliminaries: Learning interaction
We formalise learning as an ongoing interaction between a learner and an environment, evolving over discrete time steps $t\in{\mathbb{N}}_{0}=\{0,1,\dots\}$. We distinguish between two sources of stochasticity:
  • External probabilities $p_{e}(\cdot)$ describe the stochasticity inherent to the environment.
  • Predictive distributions $q_{f}(\cdot)$ describe the predictive uncertainty of the learner.
To formalise this interaction, we introduce the notion of an environment-learner interface, which provides a structure for the data exchanged between them at each time step.
Definition 3.1 (Interface).
An environment-learner interface is a pair $(\mathcal{X},\mathcal{Y})$ of measurable spaces.
Elements of $\mathcal{X}$ are observations, encompassing the signals the environment emits (e.g., features, targets, rewards, etc.), and elements of $\mathcal{Y}$ are outputs, encompassing the signals emitted by the learner (e.g., actions, classifications, regression values, etc.).
Definition 3.2 (Histories).
The set of histories relative to an interface $(\mathcal{X},\mathcal{Y})$ is
$\mathcal{H}=\bigcup_{t=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{t+1}.$ (2)
We refer to an element of $\mathcal{H}$ as a history $H\in\mathcal{H}$, which is a sequence $H_{0:t}=((X_{0},Y_{0}),\dots,(X_{t},Y_{t}))$ of $t+1$ observation-output pairs $(X_{i},Y_{i})$. For a sequence $A=(A_{0},A_{1},\dots,A_{t})$ and indices $0\leq i\leq j\leq t$, we denote the subsequence between indices $i$ and $j$ as $A_{i:j}=(A_{i},A_{i+1},\dots,A_{j})$.
Definition 3.3 (Environment).
An environment relative to interface $(\mathcal{X},\mathcal{Y})$ is a pair $(e,p_{X_{0}})$, where
  • $e:\mathcal{H},\mathcal{Y}\to\mathcal{D}(\mathcal{X})$ assigns to each history-current output pair $(H,Y)\in(\mathcal{H},\mathcal{Y})$ a conditional distribution $p_{e}(\cdot\mid H,Y)$ over the next observation,
  • $p_{X_{0}}\in\mathcal{D}(\mathcal{X})$ specifies the distribution of the initial observation $X_{0}$.
Definition 3.4 (Learner).
A learner relative to interface $(\mathcal{X},\mathcal{Y})$ is a tuple $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$, consisting of:
  • a measurable space $\mathcal{Z}$, called the learner state space;
  • a prediction function $f:\mathcal{Z}\times\mathcal{X}\to\mathcal{D}(\mathcal{Y})$, giving conditional distributions $q_{f}(\cdot\mid z,x)$;
  • a learning-mode state update function $u:\mathcal{Z}\times\mathcal{X}\times\mathcal{Y}\to\mathcal{D}(\mathcal{Z})$;
  • an inference-mode state update function $u^{\prime}:\mathcal{Z}\times\mathcal{X}\times\mathcal{Y}\to\mathcal{D}(\mathcal{Z})$;
  • an initial learner state distribution $p_{Z_{0}}\in\mathcal{D}(\mathcal{Z})$.
Although a learner’s entire state could, in principle, be updated by a single function (as in work by dong2022simple, dong2022simple and kumar2025continual, kumar2025continual), we distinguish between two update functions to capture different modes of learner evolution. During interaction with the environment, the learning-mode update $u$ governs the evolution of the entire state, including predictive parameters and auxiliary components, such as replay buffers. In contrast, the inference-mode update $u^{\prime}$ allows auxiliary components of the state, such as buffers or counters, to evolve while keeping the predictive parameters fixed. This distinction allows an observer to analyse the learner’s behaviour in both dynamic conditions, where beliefs are updated over time, and in static conditions, where beliefs remain fixed.
The interaction process.
The interaction between a learner and an environment defines a joint stochastic process over observations $X\in\mathcal{X}$, outputs $Y\in\mathcal{Y}$, and states $Z\in\mathcal{Z}$.
Definition 3.5 (Interaction Process).
The (learning-mode) interaction process between an environment $(e,p_{X_{0}})$ and a learner $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$ relative to an interface $(\mathcal{X},\mathcal{Y})$ is the stochastic process over $(X_{t},Y_{t},Z_{t})$ defined by:
Initialisation: $Z_{0}\sim p_{Z_{0}},\quad X_{0}\sim p_{X_{0}},$
Interaction: $Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),$ (learner samples output)
$X_{t}\sim p_{e}(\cdot\mid H_{0:t-1},Y_{t}),$ (environment samples current observation)
$Z_{t}\sim u(Z_{t-1},X_{t},Y_{t}).$ (learner updates state)
This interaction process generates the stochastic processes $X_{t}\,(t\geq 0),\,Y_{t}\,(t>0),\,\text{and }Z_{t}\,(t\geq 0)$.
3.2 Induced future distributions
Figure 1: State evolution and induced futures. The learner’s internal state $Z_{t}$ evolves through training updates $u$ given $(Z_{t-1},X_{t},Y_{t})$. Each state induces a distribution over induced futures $q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t})$, which is updated in inference mode $u^{\prime}$. This separation illustrates how each state encodes retained capabilities, and how training vs. introspective updates interact.
At any time during learning, the learner’s state encodes expectations about how future interactions will unfold. We can make these expectations explicit by asking the learner to internally simulate the future. Rather than waiting for new observations, we let the learner “roll forward” in inference mode using update $u^{\prime}$ to project how the sequence of inputs and outputs may continue. This is very similar to placing a learner into inference mode, where it projects how the sequence of outputs might continue.
We refer to such a simulation as an induced future. It represents a hypothetical rollout of the learner’s predictive model under its current beliefs, abstracting away interaction with the external data stream. To distinguish this hypothetical evolution from real learning progression, we introduce a futures time index $s\in\mathbb{N}$ and denote steps within the simulated trajectory as superscripts on $X$, $Y$, and $Z$.
During the rollout, the learner generates targets from its own predictive distribution, while any components not represented by the learner are borrowed from the environment. Formally, given a history $H_{0:t}$ and state $Z_{t}$, the induced future evolves according to
$Y^{s}\sim q_{f}(\cdot\mid Z_{t}^{s-1},X^{s-1}),\quad X^{s}\sim q_{e}(\cdot\mid H_{0:s-1},Y^{s}),\quad Z_{t}^{s}\sim u^{\prime}(Z_{t}^{s-1},X^{s},Y^{s}),$ (3)
where $q_{e}$ is a hybrid distribution that treats the learner’s predictions as targets while borrowing components from the environment as needed, $u^{\prime}$ is the learner’s inference mode update function, and $s=(t+1,t+2,\dots)$. Progression along the futures index $s$ provides a principled means to analyse the learner independently of new observations. We emphasise that this futures distribution is entirely isolated from the interaction process.
We refer to an element of future histories as $H\in\mathcal{H}$, which is a sequence of future observation-output pairs $(X^{i},Y^{i})$ induced by the learner $H^{t+1:\infty}=((X^{t+1},Y^{t+1}),(X^{t+2},Y^{t+2}),\dots)$.
Definition 3.6 (Induced Future Distributions).
The induced future distribution of the learner at state $Z_{t}$ with realised history $H_{0:t}$ is the joint distribution
$q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ (4)
defined by rolling out the inference-mode process in (3).
Thus, an induced future is a probability distribution over entire infinite sequences of future inputs and outputs, i.e., elements of $\mathcal{X}\times\mathcal{Y}$, which we denote by $(\mathcal{X}\times\mathcal{Y})^{\mathbb{N}}$. Because the learner’s state $Z_{t}$ and subsequent observations are random, the induced future distribution at each step $q(\cdot\mid Z_{t},H_{0:t})$ is a random variable taking values in $\mathcal{D}((\mathcal{X}\times\mathcal{Y})^{\mathbb{N}})$.
Predictive-Bayesian perspective.
The construction above motivates a predictive perspective of learning, conceptually inspired by the predictive Bayesian framework (fortini2019quasi; fortini2025exchangeability; fong2023martingale). In this formulation, the learner’s state $Z_{t}$ is characterised by its predictive distribution over future observations rather than by latent parameters or priors. This distinction is crucial in deep learning, where network parameters serve primarily as computational intermediaries and often lack interpretable statistical meaning. By tracking the evolution of predictions instead of parameter updates, we obtain a formulation of learning that remains empirically interpretable.
A further advantage of this predictive perspective is that it can be validated: predictive statements can be evaluated against realised outcomes, providing an immediate criterion for model adequacy (fong2023martingale). In contrast, probability statements about latent variables, such as priors or posteriors, do not admit direct validation. Framing beliefs in predictive terms also allows for discussion of stationarity with respect to the learner: whether a learner’s sequential forecasts remain self-consistent as learning progresses. This notion of predictive self-consistency provides the conceptual foundation for analysing the dynamics of learning and, ultimately, for our definition of forgetting in § 4.2.
3.3 Learning as a stochastic process
The stochasticity of the environment induces stochasticity in the learner’s state evolution; as the learner interacts with the environment, the sequence $\{Z_{t}\}$ forms a stochastic process. Since each state $Z_{t}$ defines an induced future distribution $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$, the sequence of induced future distributions over time also forms a stochastic process. Equivalently, the distribution over learner states at time $t$ defines a distribution over distributions of futures, capturing the learner’s evolving uncertainty about possible continuations. This is illustrated in Figure 1.
Takeaway 3.1 The stochastic sequence of observations drives a stochastic evolution of states $\{Z_{t}\}$, which in turn induces a stochastic evolution of induced future distributions $\{q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\}$.
Interpretation across learning paradigms.
The abstract variables $(X_{t},Y_{t},Z_{t})$ admit natural interpretations across learning settings. In supervised learning, $X_{t}$ denotes (current input, previous target) pairs and $Y_{t}$ the learner’s output in response to the previous input; in RL, $X_{t}$ contains states and rewards while $Y_{t}$ are actions; in generative modelling, $Y_{t}$ directly models $X_{t}$.
The learner state $Z_{t}$ encapsulates all contents of the learner (including parameters, latent variables, buffers, etc.). Distinct modes of operation correspond to different update rules: in training (update rule $u$), both the learner’s belief and auxiliary structures may adapt; in inference ($u^{\prime}$), the learner’s belief remains fixed while transient components continue to evolve.
For example, in a supervised classification setting, $X_{t}$ consists of a (current input, previous target) pair, $Y_{t}$ is the label predicted by the learner given $X_{t-1}$, and $Z_{t}$ may represent parameters and momenta. The learning-mode update $u$ performs gradient steps on a loss that compares the observed and predicted labels. The inference-mode update $u^{\prime}$ does not change the parameters and momenta.
Thus, while the same symbols are used throughout, their interpretation shifts depending on the paradigm. What remains consistent is the structural principle: the learner generates $Y_{t}$ conditioned on $(Z_{t-1},X_{t-1})$, the environment supplies the observation $X_{t}$, and the state $Z_{t}$ evolves via $u$. This defines a single stochastic-process formalism that encompasses general learning processes.
4 Forgetting
We consider an observer monitoring the interaction between a learner and its environment. The observer evaluates how the learner’s capabilities evolve over time, with particular emphasis on quantifying the extent to which the learner forgets during the interaction.
4.1 Characterising forgetting
Forgetting is the loss of previously acquired information or capabilities (jagielski2022measuring). To formalise this notion, we first outline desiderata that any valid measure should satisfy, guided by thought experiments (see Appendix C for detailed scenarios).
Desideratum 1.
A forgetting measure should quantify the loss of learned information over time.
Forgetting is distinct from task performance; a learner may maintain outdated or incorrect beliefs yet still perform well, or may discard relevant knowledge without an immediate drop in performance. A good measure of forgetting must therefore disentangle information loss from performance fluctuations.
Desideratum 2.
A characterisation of forgetting must not conflate forgetting with the correctness of outputs or with justified updates that change beliefs.
For instance, when a learner revises its predictions in light of new observations, its state changes. This update is not necessarily forgetting, since this state change can occur without any loss of information.
Desideratum 3.
A forgetting measure must capture losses in a learner’s prior capabilities, not just the retention of previously observed data.
Forgetting includes the loss of general capabilities, not just individual observations. For instance, a learner may fail to generalise to unobserved examples that it previously handled correctly. Measures of forgetting that only focus on memorisation overlook this broader notion.
Desideratum 4.
Forgetting should characterise the learner’s loss of information and capabilities resulting from the learner’s adaptations.
Forgetting is a property of the learner, not a property of the environment or dataset in which it operates. For example, a system that never adapts cannot meaningfully forget. Therefore, a learner’s environment may influence the rate or magnitude of forgetting, but it does not define it.
These desiderata provide a principled test bed for developing and evaluating measures of forgetting. The complete set of thought experiments in Appendix C illustrates the need for these desiderata and motivates the formalism developed below.
4.2 Forgetting
Learning is inherently a process of both absorption and loss. Each new observation provides the learner with new information while simultaneously displacing or overwriting previously learned information (as illustrated in Figure 2). This interdependence is fundamental: every update reshapes what the learner can represent and, consequently, what it may forget.
Figure 2: Information absorption and loss over a single learning update. The figure illustrates the information dynamics occurring within each learning step. At each update, the previous state, $Z_{t-1}$, receives a new observation, $X_{t}$. The learner incorporates information from $X_{t}$ into its state while simultaneously displacing part of the information previously encoded in $Z_{t-1}$. This process results in a new state $Z_{t}$, reflecting the assimilation of new information and the loss of some prior information.
If a learner’s behaviour changes when exposed to data it already expects, it can only be losing information, since no new information has been introduced. The learner’s expectations are defined by its predictive distribution for a given observation. We leverage this intuition to formalise forgetting.
Recall that the learner maintains a state that evolves recursively:
$Z_{t}\sim u(Z_{t-1},X_{t},Y_{t}).$ (5)
At each time $t$, the state $Z_{t}$ induces a distribution over futures $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$, which characterises both the learner’s current predictive capabilities and its expectations about futures.
We therefore define forgetting in terms of induced futures, rather than the learner’s state or predictions. Different states–whether parametric, non-parametric, or otherwise–may induce identical futures:
$q(H^{t+1:\infty}\mid Z_{t},H_{0:t})=q(H^{t+1:\infty}\mid Z_{t}^{\prime},H_{0:t}),$ (6)
where $Z_{t}\neq Z_{t}^{\prime}$. Furthermore, predictions about already-seen observations may shift without indicating that forgetting has occurred. By grounding our definition in induced futures, we ensure that updates are evaluated with respect to the learner’s expected predictions, thereby distinguishing between constructive and destructive adaptation. This satisfies 2 and 3.
Consistency as non-forgetting.
Building on this intuition, a learner is unforgetful if updating on targets that are entirely consistent with its own predictions does not change its induced futures. The distribution of futures is expected to remain invariant when no new information is provided to the learner. Notably, suppose a learner’s update function $u(Z_{t-1},X_{t},Y_{t})$ depends on the history $H_{0:t-1}$. In that case, consistency may be violated and forgetting can occur (see § B.3 for a detailed discussion).
Formally, let $q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})$ denote the distribution over futures before updates on targets consistent with the learner’s expectations, and let $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ denote the distribution after updating on targets consistent with the learner’s expectations. Non-forgetting requires:
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t},X_{t},Z_{t}}\left[q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\right],$ (7)
where $Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1})$, $X_{t}\sim q_{e}(\cdot\mid H_{0:t-1},Y_{t})$, and $Z_{t}\sim u(\cdot\mid Z_{t-1},X_{t},Y_{t})$.
In these updates, $X_{t}$ is sampled from the same hybrid distribution $q_{e}$ introduced in § 3.2. Updates are performed on these learner-consistent targets to ensure a separation between forgetting and backward transfer. In this formulation, the expectation in (7) is taken over both the stochasticity of the learner’s target generation and the environmental inputs, yielding a description of predictive consistency.
This notion extends naturally to multiple updates. Over $k$ steps, non-forgetting requires that futures induced before any updates remain compatible with those induced after $k$ interactions have occurred. Equivalently, earlier futures must be exactly recoverable by marginalising over all possible $k$-step interaction paths, leading to the generalised consistency condition.
Definition 4.1 (Consistency Condition).
For $k\geq 1$, a learner is $k$-step consistent if and only if
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{X_{t:t^{\prime}},Y_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{0:t^{\prime}})\right],$ (8)
where $t^{\prime}=t+k-1$, and for $i=t,\ldots,t^{\prime}$ the expectation is taken over
$X_{i}\sim q_{e}(\cdot\mid H_{0:i-1},Y_{i}),\quad Y_{i}\sim q_{f}(\cdot\mid Z_{i-1},X_{i-1}),\quad Z_{i}\sim u(\cdot\mid Z_{i-1},X_{i},Y_{i}).$
Propensity to forget.
In practice, most learning algorithms do forget. To quantify how much the learner is likely to forget at any time $t$, we introduce a concrete measure grounded in our formalism. When the consistency condition is violated, the learner’s updated induced futures diverge from its initial induced futures. Measuring this divergence yields a natural notion of a learner’s propensity to forget. This operationalises our conceptual definition, allowing us to validate the formalism by ensuring that the measure aligns with intuitive expectations about forgetting in practice.
Definition 4.2 (Propensity to Forget).
The $k$-step propensity to forget incurred at time $t$ is measured as the divergence
$\Gamma_{k}(t)\coloneqq\mathrm{D}\left(q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\,\|\,q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\right)$ (9)
where $\mathrm{D}(\cdot\|\cdot)$ is a suitable divergence measure.
Takeaway 4.1 Forgetting occurs when the consistency condition is violated; the induced futures after $k$ updates are no longer recoverable from those before the updates.
Scope and boundary of validity.
The perspective of the predictive Bayesian formalism assumes that the learner’s internal state parameterises a coherent predictive distribution over futures: that is, there exists a well-defined mapping
$Z_{t}\mapsto q(H^{t+1:\infty}\mid Z_{t},H_{0:t}),$ (10)
such that all of the learner’s expectations about future interactions arise from a single, self-consistent probabilistic model. This condition ensures that the learner’s predictive state fully captures its beliefs about the future and that updates to these beliefs evolve within a single probability space. When this mapping exists, the induced futures $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ are well-defined and measurable, allowing the propensity to forget to be meaningfully evaluated. In practice, coherence is typically achieved throughout training, when the learner’s state defines a consistent model of its futures. During certain transitory phases, however, such as buffer reinitialisation, target-network lag, or policy resets, this mapping may temporarily cease to exist. In those intervals, the learner’s state no longer induces a coherent distribution over futures, and the notion of forgetting becomes undefined, not because the formalism fails, but because the learner itself momentarily lacks a predictive model of its behaviour.
More generally, some algorithms may never instantiate such a coherent predictive mapping during their operation. These systems, by construction, do not admit a single probabilistic interpretation of their evolving state and thus fall outside the scope of this formalism. The previously identified conditions of failure (when the learner’s updates do not correspond to transformations within a unified predictive model family) delineate this boundary. In all other cases, the induced futures remain well-defined whenever coherence is achieved, and the forgetting formalism applies without modification.
5 Empirical analysis
We now examine forgetting empirically. Our theoretical account provides a general conceptualisation of forgetting. To illustrate its utility, we instantiate and study Definition 4.2 across multiple environments and learning algorithms.
Figure 3: Forgetting occurs across all deep learning scenarios. Left: Forgetfulness dynamics of a shallow neural network trained on regression, classification, and generative modelling tasks. The solid line shows the $k$-step forgetfulness (where $k$ varies from 11 to 4040) over the normalised training step. Regression and classification tasks use KL divergence, while the generative task uses the maximum mean discrepancy (MMD). Forgetting dynamics vary throughout training, even without any distribution shift. Right: A class-incremental learning example using a single-layer neural network on a two-moons classification task. We show the $k$-step forgetting profile over four seeds, with the shaded area indicating the spread of $\Gamma_{k}(t)$ over $k$ from 1 to 40. The plot illustrates the abrupt increase in forgetting at the task boundary. See Appendix F for details on the experimental implementation.
5.1 Forgetting profiles in different settings
Across all examined settings, forgetting is non-zero, and both its dynamics and magnitude vary significantly (see Figure 3): this motivates our paper title, “Forgetting is Everywhere”. Absolute $\Gamma_{k}(t)$ values are domain-specific, but the shape and dynamics of the forgetting trajectory convey meaningful information. As shown in Figure 4, even minor changes to the learning algorithm can substantially alter $\Gamma_{k}(t)$.
5.2 Approximate learners can benefit from forgetting
At each update, an approximation-based learner incorporates new information from current observations while discarding parts of its existing state (§ 4.2). Because approximate updates yield imperfect representations, a learner’s performance depends on striking a balance between adapting to new information and retaining useful prior information.
To study this effect, we investigate how modifications to the learner influence the propensity to forget. Across experiments, a consistent pattern emerges for approximate learners: a moderate amount of forgetting improves learning efficiency. Here, we quantify training efficiency using the inverse of the normalised area under the training loss curve, a practical proxy for learning speed and convergence quality. Empirically, the forgetting-efficiency relationship shows an “elbow” (Figure 4), indicating that optimal training efficiency occurs at a non-zero level of forgetting. This suggests that effective approximate learners utilise forgetting as a mechanism for adaptive and efficient learning.
Takeaway 5.1 Forgetting is ubiquitous in deep learning. The trade-off between training efficiency and forgetting determines the optimal amount to forget–in deep learning, this is rarely zero.
5.3 Implications of distribution shift
Distribution shift and stochasticity strongly influence forgetting dynamics. In i.i.d. environments, the interaction process is stationary and induced futures evolve smoothly. When learner hyperparameters are well-tuned and training conditions are stable, forgetting dynamics become smooth and predictable, allowing the learner to balance adaptation and retention effectively. In such settings, training efficiency is high because the learner’s updates operate under consistent conditions, and the underlying data distribution provides consistent feedback for improvement.
Figure 4: Approximate learners can benefit from non-zero forgetting. We analyse how training efficiency and forgetfulness co-vary across learning algorithms. Training efficiency is measured as the inverse of the normalised area under the training loss curve, an approximate but informative proxy for learning speed and convergence quality in the setting we study. Forgetfulness is quantified as the mean 40-step propensity to forget, $\bar{\Gamma}_{40}(t)$, across training steps. Left: In a regression task, varying stochastic gradient descent’s momentum parameter shows that higher momentum increases forgetfulness, with maximum training efficiency at $0.9$ momentum. Right: Varying model size shows that maximum efficiency occurs at 20 parameters. In both cases, the most efficient learners exhibit some forgetting–too little slows adaptation, too much destabilises learning–highlighting a fundamental trade-off.
In CL, abrupt shifts in the observation distribution cause discontinuous changes in $Z_{t}$ and its induced futures. Consequently, the magnitude of consistency violation abruptly increases at task boundaries (Figure 3) as the learner must rapidly adapt to a new task.
RL, however, remains non-stationary throughout the training process. The learner’s data distribution evolves in response to its policy, creating feedback loops that continually perturb its state. This interplay between environmental and algorithmic non-stationarity induces chaotic fluctuations in the forgetting profile (Figure 5). Current neural network optimisation algorithms, designed for the more stationary supervised learning setting, struggle to stabilise these dynamics. As a result, RL learners often fail to maintain a consistent balance between retention and adaptation, leading to poor learning dynamics and reduced training efficiency.
Takeaway 5.2 Stable forgetting dynamics improve learning efficiency by promoting a balanced trade-off between adaptation and retention.
Figure 5: Chaotic forgetting dynamics in RL. We show TD loss and Q-value evaluation across ten seeds for a DQN learner trained on the cartpole environment. While neural networks trained on supervised learning can exhibit fluctuating forgetting dynamics, when hyperparameters are well-tuned, their dynamics are typically smooth and self-stabilising. In contrast, neural networks trained via RL display persistent, chaotic fluctuations driven by non-stationarity in the environment. Current RL algorithms poorly regulate these dynamics, resulting in unstable learning and reduced efficiency.
6 Conclusion
In this work, we have provided a conceptual foundation for understanding forgetting in general learning systems from a predictive perspective. We showed that forgetting can be understood as a violation of self-consistency in a learner’s predictive distribution and formalised the conditions under which forgetting occurs. We proposed an operational measure, the learner’s propensity to forget (Definition 4.2), which allowed us to validate this formalism in practice. Our results demonstrate that measures derived from the predictive formulation faithfully capture the extent of forgetting in practice and provide a clear distinction between constructive and destructive adaptation.
Empirically, we found that forgetting is pervasive across diverse settings and is determined by the interaction between the learner and the environment. Furthermore, we identified a trade-off between forgetting and training efficiency, with CL and RL settings particularly susceptible to forgetting and thus exhibiting poorer training efficiency than other learning settings.
By providing a conceptual formalism of forgetting, our work offers a principled foundation for analysing and understanding forgetting. We hope this perspective will guide the development of learning algorithms that can adapt effectively without sacrificing prior abilities, and provide a better conceptual basis for reasoning about how a learner’s capabilities emerge, persist, and degrade over time.
Acknowledgements
The authors would like to thank Guy Frankel, Jonathon Hare, Henry Gouk, Andras Szecsenyi and Matej Sandor for their help and valuable insights. This research was enabled by computer resources provided by the School of Informatics at the University of Edinburgh. Funding for this research was provided by EPSRC through a PhD studentship within the CDT in Machine Learning Systems hosted in the School of Informatics, University of Edinburgh (EP/Y03516X/1).
Appendix
The appendix is organised as follows: Appendix A defines the notation, Appendix B expands on the theory, Appendix C gives illustrative thought experiments, Appendix D provides additional details regarding empirical implementation, Appendix E reports additional results, and Appendix F details our experiments.
–appendices continue on next page–
Appendix A Notation
We first provide a table summarising the notation used throughout the paper.
Table 1: A summary of notation.
Symbol Description Domain / Type
Interface-related
$(\mathcal{X},\mathcal{Y})$ Environment–learner interface Pair of measurable spaces
$\mathcal{X}$ Observation space Measurable space
$\mathcal{Y}$ Prediction space Measurable space
$\mathcal{H}$ Set of histories $\bigcup_{t=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{t}$
$X_{t}$ Observation at time $t$ Element of $\mathcal{X}$
$Y_{t}$ Output at time $t$ Element of $\mathcal{Y}$
$H_{0:t}$ History up to time $t$ Sequence $((X_{1},Y_{1}),\dots,(X_{t},Y_{t}))$
Environment-related
$(e,p_{X_{1}})$ Environment $p_{X_{1}}\in\mathcal{D}(\mathcal{X})$,  $e:\mathcal{H}\to\mathcal{D}(\mathcal{X})$
$p_{e}(\cdot\mid h,y)$ Environment distribution given history $h$ and output $y$ $\mathcal{D}(\mathcal{X})$
$q_{e}(\cdot\mid h,y)$ Learner-sampled hybrid input distribution given history $h$ and output $y$ $\mathcal{D}(\mathcal{X})$
Learner-related
$(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$ Learner Tuple
$\mathcal{Z}$ Learner state space Space
$Z_{t}$ Learner state at time $t$ Element of $\mathcal{Z}$
$Z_{t}^{s}$ Learner state in inference mode at a futures time $s$ Element of $\mathcal{Z}$
$p_{Z_{0}}$ Initial learner state distribution $\mathcal{D}(\mathcal{Z})$
$f$ Prediction function $\mathcal{Z}\times\mathcal{X}\to\mathcal{D}(\mathcal{Y})$
$u$ Learning-mode state update $\mathcal{Z}\times\mathcal{X}\times\mathcal{Y}\to\mathcal{D}(\mathcal{Z})$
$u^{\prime}$ Inference-mode state update $\mathcal{Z}\times\mathcal{X}\times\mathcal{Y}\to\mathcal{D}(\mathcal{Z})$
$q_{f}(\cdot\mid z,x)$ Predictive distribution $\mathcal{D}(\mathcal{Y})$
Induced futures
$q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ Induced future distribution $\mathcal{D}((\mathcal{X}\times\mathcal{Y})^{\mathbb{N}})$
$X^{s}$ Future observation at future time $s$ Element of $\mathcal{X}$
$Y^{s}$ Future output at future time $s$ Element of $\mathcal{Y}$
$H^{t+1:\infty}$ Sampled induced future at time $t$ Sequence in $((X^{t+1},Y^{t+1}),(X^{t+2},Y^{t+2}),\dots)$
A.1 Framework generality
We now provide explicit examples of different realisations of the abstract variables introduced in § 3.
Table 2: This table provides specific examples of values for the abstract variables: $X_{t},Y_{t}$ and $Z_{t}$ across different common machine learning settings.
Setting $X_{t}$ (observation) $Y_{t}$ (prediction) $Z_{t}$ (learner state)
Regression Current input, previous target pair $(x_{t},\;y_{t-1})$ Predicted target $\hat{y}_{t-1}$ Parameters (e.g., weights, optimiser state)
Classification Current input, previous label pair $(x_{t},\;\ell_{t-1})$ Predicted class label $\hat{\ell}_{t-1}$ Classifier parameters, latent state (e.g., neural network weights, momenta)
Generative modelling Data to generate $x_{t}$ Attempted generations of $x_{t}$ Generative model parameters, latent variables, buffers
Bayesian models Data (possibly $(x_{t},y_{t-1})$ pairs) Posterior predictive sample $\hat{y}_{t-1}$ Posterior / variational parameters, latent variables
Reinforcement learning State–reward $(s,r)$ or transition $(s,a,r,s^{\prime})$ Action $a$ (policy output) or predictive model of next state Policy/value parameters, transition/reward models, replay buffer, auxiliary structures
–appendices continue on next page–
Appendix B Theoretical results
Recall, the learner’s state is updated according to
$Z_{t}\sim u(Z_{t-1},X_{t},Y_{t}).$ (11)
Before observing $X_{t},Y_{t}$ and updating to $Z_{t}$, we have a futures distribution
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1}).$ (12)
B.1 One-step consistency condition
After the update, a non-forgetting learner satisfies the following conditional independence rules:
[…]
Furthermore, the environment admits the following independence:
[…]
The one-step consistency operator is obtained by marginalising over observations $X_{t},Y_{t}$ and updated state $Z_{t}$:
$q^{*}(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\int q(H^{t+1:\infty},X_{t},Y_{t},Z_{t}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$ (15)
This can factorised, with the conditional independence rules in (13) and (14) applied, to derive the following:
$\int q(X_{t},Y_{t},Z_{t},H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}$ (16)
$=\int q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\,u(Z_{t}\mid Z_{t-1},X_{t},Y_{t})$
$\qquad\qquad q_{e}(X_{t}\mid H_{0:t-1},Y_{t})\,q_{f}(Y_{t}\mid Z_{t-1},X_{t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$
which is the 1-step consistency condition from (7):
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t},X_{t},Z_{t}}\left[q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\right],$ (17)
where
$Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),\quad X_{t}\sim q_{e}(\cdot\mid H_{0:t-1},Y_{t}),\quad Z_{t}\sim u(\cdot\mid Z_{t-1},X_{t},Y_{t}).$
B.2 k-step consistency condition
We can apply the same reasoning to derive the $k$-step consistency condition. Denoting $t^{\prime}$ as $t^{\prime}=t+k-1$, we get:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ (18)
$=\int q(X_{t:t^{\prime}},Y_{t:t^{\prime}},Z_{t:t^{\prime}},H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t:t^{\prime}}\,\mathrm{d}Y_{t:t^{\prime}}\,\mathrm{d}Z_{t:t^{\prime}}$
$=\int q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{0:t^{\prime}})\cdot\prod_{s=t}^{t^{\prime}}\big[u(Z_{s}\mid Z_{s-1},X_{s},Y_{s})$
$\qquad q_{e}(X_{s}\mid H_{0:s-1},Y_{s})q_{f}(Y_{s}\mid Z_{s-1},X_{s-1})\big]\mathrm{d}X_{t:t^{\prime}}\,\mathrm{d}Y_{t:t^{\prime}}\,\mathrm{d}Z_{t:t^{\prime}},$
which is the $k$-step consistency condition from Definition 4.1:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t:t^{\prime}},X_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{t^{\prime}})\right].$ (19)
where, for $i=t,\dots,t^{\prime}$
$Y_{i}\sim q_{f}(\cdot\mid Z_{i-1},X_{i-1}),\quad X_{i}\sim q_{e}(\cdot\mid H_{0:i-1},Y_{i}),\quad Z_{i}\sim u(\cdot\mid Z_{i-1},X_{i},Y_{i}).$
B.3 The theoretical need for replay
The above derivations make explicit the role of the learner’s state update function $u(Z_{t-1},X_{t},Y_{t})$ in ensuring the consistency of the induced futures. In particular, the one-step and $k$-step consistency conditions rely on the assumption that the update function is such that $Z_{t}$ is conditionally independent of the whole history $H_{0:t-1}$ given $(Z_{t-1},X_{t},Y_{t})$.
However, in practice, many learning algorithms violate this condition of conditional independence, and the update function explicitly depends on the history. When this dependence exists, correctly performing consistent updates depends on access to past data. Replay mechanisms provide an empirical solution: by placing past observations into the current observation, the learner effectively restores the conditional independence assumed by the theoretical consistency conditions.
Thus, from a theoretical perspective, the need for replay arises naturally whenever the state update $u(Z_{t-1},X_{t},Y_{t})$ is insufficient at capturing all historical dependencies. Replay promotes consistent induced futures by explicitly placing much of the history into the current observation and, as a consequence, prevents forgetting that would otherwise occur.
–appendices continue on next page–
Appendix C Thought experiments
We present a series of thought experiments designed to stress-test potential definitions of forgetting in general learning systems. Each scenario illustrates a conceptual edge case that any robust definition of forgetting should naturally handle. For each scenario, we provide a consistency verdict, indicating whether our framework classifies the learner to be forgetting or not, and explain the intuition behind this judgement.
Scenario C.1 (The degenerate learner.).
Consider a learner that never updates its beliefs–its state is fixed for all time. Such a learner never changes its state or acquires new information and is thus degenerate. Does this learner ever forget? One might argue that forgetting is defined relative to the information available to the learner over time; however, if nothing is ever learned, how can anything ever be forgotten?
Consistency verdict: no forgetting. Since the state remains constant across any number of updates, the induced futures distributions are identical. A learner that never changes its state can never forget.
Scenario C.2 (0-bit and 5-bit stacks.).
Consider two first-in, first-out (FIFO) stacks: one with capacity 0 bits, another with 5 bits. The 0-bit stack cannot store information; the 5-bit stack can.
Consistency verdict: 0-bit does not forget, the 5-bit stack does. The predictive distribution of the stack is always the returned bit of the address being accessed at each step. A 0-bit stack’s state never changes, so its future remains constant. The 5-bit stack, however, overwrites past bits when it is full; its predictive distribution changes as bits are dropped, hence it forgets. Forgetting only exists in systems capable of learning.
Scenario C.3 (The hash map.).
A hash map with infinite associative memory never overwrites existing entries. It represents a theoretical “perfect rememberer”.
Consistency verdict: no forgetting. The predictive distributions of such a learner are a Dirac at the retrieved value (or null if unassigned), and the induced futures remain identical over time. Thus, the learner never forgets.
Scenario C.4 (The clock.).
Consider a simple clock that increments its state deterministically with each tick. Does the clock forget past states? In one sense, yes: it overwrites its register and never recovers earlier times. In another sense, no: it never misrepresents the current time and updates consistently. This raises the question: must forgetting be tied to the loss of recoverable information, or only to deviations from self-consistent updates?
Consistency verdict: no forgetting. Under regular environmental updates, the clock advances its state (its notion of “time”) by one unit per tick. However, under consistency updates, we propagate the learner’s state forward using its own current beliefs rather than the actual environmental progression. Since the clock believes time remains fixed at its last known state, it will continue to predict the exact time indefinitely. Consequently, its futures remain invariant under consistency updates. The clock does not forget.
Scenario C.5 (The moody learner.).
Suppose a learner only updates on even timesteps, ignoring all odd ones. Does inaction indicate perfect memory?
Consistency verdict: no forgetting (on ignored steps). When no update occurs, the state remains identical, so induced futures distributions are unchanged. Forgetting can only occur during state-altering updates.
Scenario C.6 (The function picker.).
Suppose there are $L$ learner functions, and at each timestep, one function is selected uniformly at random, independent of both $t$ and the observed data.
Consistency verdict: no forgetting. In expectation, the induced futures after any number of consistent updates remain invariant. The randomness of selection does not imply loss of information about possible futures.
Scenario C.7 (The binary flipper.).
Consider a binary classification model that flips its predictions at every timestep.
Consistency verdict: no forgetting. Similar to the clock, the consistency update believes time remains fixed at its last known state; thus, the learner does not flip its predictions, and its state remains fixed. Consequently, its futures remain invariant under consistency updates, and the flipper does not forget.
Scenario C.8 (Label permutation.).
Take a perfectly trained binary classifier. Now, consider that we permute the mapping between logits and output labels without changing the parameters. On the one hand, the parameters remain the same; on the other, the outputs have changed significantly. Has such a model forgotten everything it knew just because its behaviour has changed, or has it not forgotten anything at all because it has the same states?
Consistency verdict: no forgetting. Similar to the clock, the consistency update believes time remains fixed at its last known state; thus, the learner does not flip its predictions during consistent updates, and its state and predictive distributions remain fixed. Consequently, its futures remain invariant under consistency updates, and the learner does not forget.
Scenario C.9 (Forgetting unseen but generalised inputs.).
Consider a learner that is trained on the MNIST dataset. At time $t$, the model correctly classifies all instances of the digit “4” in the test set. After further training, it loses the ability to correctly classify some test examples. Can a model forget something that it has never directly encountered (such as specific test inputs) but has nonetheless learned to generalise to?
Consistency verdict: this forgets. Even though the test data were unseen, the learner’s induced futures change when its performance on those data changes. Thus, the learner has forgotten: forgetting occurs whenever the previously supported predictive capabilities vanish, regardless of whether the corresponding data have been observed.
Scenario C.10 (Even number checkers.).
Suppose a model receives binary inputs sequentially and deterministically predicts a $1$ if it has observed an even number of $1$s, and 0 otherwise. Does it forget because its state changes over time?
Consistency verdict: no forgetting. The next $k$ observations may either be a 0 or a $1$, drawn from any distribution. After $k$ updates, the learner’s predictive distribution is deterministically either a 0 or a $1$. Although the state continues to evolve across the induced futures, the expected predictive distribution remains the same across them. Therefore, no forgetting occurs.
Scenario C.11 (Surprising events.).
Consider the scenario where a fair coin is flipped repeatedly. After 10 coin flips, we are in an unlikely situation where all flips have so far resulted in heads. If, on the next flip, we observe a tail, the learner may be surprised, even if it has remembered all previous results.
Consistency verdict: no forgetting. Under consistent updates, future updates are performed relative to the learner’s current beliefs rather than the actual environment. This distinction is what prevents surprise from being mistaken for forgetting. Since the learner’s predictive distribution already encodes its own uncertainty over future outcomes, an unexpected tail does not constitute loss of information; it is a valid observation under the learner’s model. Consequently, the expected futures distribution before and after the $k$ updates remains identical, and no forgetting is detected.
Scenario C.12 (Bayesian optimisation.).
Consider a Bayesian optimiser that initially observes high objective values, then later finds a lower minimum. Its predictive distribution must update in response to the new observation. However, no previously acquired evidence has been discarded. This is a knowledge update, not forgetting; naive comparisons of states may mistakenly classify this update as forgetting.
Consistency verdict: no forgetting. As in C.11, future updates are computed relative to the learner’s current predictive distribution rather than the actual environment. Because updates reflect the learner’s own beliefs, encountering a new minimum does not imply a loss of prior information. In expectation, the futures distributions before and after $k$ updates remain identical, indicating that the learner has not forgotten.
–appendices continue on next page–
Appendix D Empirical considerations
In this section, we detail the approach used to estimate the forgetfulness measure $\Gamma_{k}(t)$. $\Gamma_{k}(t)$ captures the extent to which the learner’s induced futures at time $t$ remain consistent with those obtained after $k$ simulated updates. While this quantity cannot be computed exactly, it can be approximated using a particle-based Monte Carlo scheme. At a given time, we clone the learner into $M$ replicas and propagate each replica forward under $k$ steps of simulated interaction with its own induced futures. This yields a set of post-update models $\{Z_{t+k-1}^{(m)}\}_{m=1}^{M}$, whose mixture distribution provides an approximation to the consistency condition:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\approx\frac{1}{M}\sum_{m=1}^{M}q(H^{t+k:\infty}\mid Z_{t+k-1}^{(m)},H^{(m)}_{t+k-1})$, (20)
We compare the learner’s distribution on a reference trajectory of simulated futures with this mixture distribution. The divergence between the two provides an empirical estimate of $\Gamma_{k}(t)$.
Predictive distributions.
In practice, the predictive distribution $q_{f}(Y_{t}\mid Z_{t-1},X_{t-1})$ is obtained directly from the model’s output parameterisation. For classification tasks, this distribution is explicit: the network’s logits define a categorical (or Bernoulli) distribution over discrete classes. A similar interpretation applies to generative models, where the predictive distribution corresponds to the model’s generative distribution over data samples. For regression tasks, the predictive distribution is determined by the likelihood implied by the training loss. When optimising a mean-squared-error objective, for example, the corresponding likelihood is Gaussian, so the model’s prediction corresponds to the mean of a normal distribution. The predictive variance can be estimated empirically from the residual error distribution on a held-out validation set. Thus, even though most practical neural networks output only point predictions, their underlying training objective defines an implicit predictive distribution. We use this to evaluate the induced futures and empirically quantify forgetting.
Visualising the propensity to forget
Figure 6 compares a single-layer neural network’s futures distribution at the current training step with the same model but after $40$ additional updates on futures data. This allows us to visualise the divergence between the two induced futures distributions. This difference quantifies the model’s propensity to forget, as defined in Definition 4.2.
Figure 6: Comparison of initial and 40-step induced futures. Axes show the input space around the two-moon binary classification task. Example datapoints are overlaid and coloured by target class ($\text{red}=0$, $\text{blue}=1$). The background grid is shaded according to the classifier’s logit values, from red (0) to blue (1), with white indicating uncertainty at 0.5. Left: The initial induced futures distribution $q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ serves as the baseline for consistent updates. Right: The $k=40$ induced futures distribution $q^{*}_{k}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$, showing how the predictive model is likely to evolve over 40 steps. The divergence in Definition 4.2 is computed between these two distributions.
–appendices continue on next page–
Appendix E Additional results
In this section, we discuss additional results, including Bayesian learners, permutation sensitivity, and an expanded discussion of § 5.
E.1 Unforgetful learners
Figure 7: Self-consistent learners do not forget. Axes showing observations $(x,y)$ are shaded white, and parameter axes $(w_{0},w_{1})$ are shaded grey. Top row: The same four observations are presented to a linear regression learner in different orders. Second row: A full Bayesian posterior represents a complete summary of all observations and satisfies the $k$-step consistency condition, implying that no forgetting has occurred. The permutation invariance of the learner in i.i.d. settings is a corollary of this property, due to Bayesian commutative updates. Bottom rows: Two constrained learners – a Gaussian variational posterior with diagonal covariance and a point estimate updated by gradient descent – violate self-consistency. Their updates alter induced futures; therefore, previously supported capabilities may be lost. The final column shows the resulting posterior predictive distributions.
We illustrate forgetting behaviour using Bayesian linear regression, where the exact Bayesian learner is analytically tractable, self-consistent, and therefore does not forget.
At time $t$, the learners beliefs are captured by its predictive distribution $p(X_{t+1:\infty}\mid\theta_{t},X_{0:t})$. Let $X_{1:t}$ denote the observations to time $t$, and let $\theta$ denote the parameters. The Bayesian posterior after observing $X_{1:t}$ is
$p(\theta\mid X_{1:t})\propto p(\theta)\prod_{i=1}^{n}p(X_{i}\mid\theta).$ (21)
Since multiplication commutes, the posterior is invariant to the order of observations:
$p(\theta\mid X_{1:t})=p(\theta\mid X_{\pi(1)},\dots,X_{\pi(t)})\quad\forall\text{ permutations }\pi.$ (22)
Therefore, exact Bayesian updates are permutation-invariant in i.i.d. settings.
Self-consistency of Bayesian updates.
Exact Bayesian learning forms a projective family. Marginalising the posterior after a hypothetical future observation recovers the current posterior:
$p(\theta\mid X_{1:t})=\int p(\theta\mid X_{1:t+1})p(X_{t+1}\mid X_{1:t})\,\mathrm{d}X_{t+1}.$ (23)
Equivalently, conditioning on a future observation and then marginalising over its prior predictive distribution returns the same belief as not conditioning at all. Thus, conditioning and marginalising commute. This is equivalent to the $k$-step self-consistency.
Self-consistency ensures that every predictive capability supported by $p(\theta\mid X_{1:t})$ remains supported after a Bayesian update unless contradicted by data. The learner preserves its prior predictive beliefs in expectation; therefore, the consistency condition is never violated, and no forgetting occurs.
Approximate procedures such as diagonal Gaussian variational inference or gradient-based point estimates do not satisfy the self-consistency condition. Their updates depend on the order of observations and yield posteriors that cannot be recovered by marginalising futures. As a result, previously supported predictions may be lost, indicating that forgetting has occurred. The predictive and parameter distributions for all learners are shown in Figure 7.
E.2 Impact of changes to algorithms
Here we expand on the empirical results presented in §5, examining how the dynamics of forgetting vary with model architecture, hyperparameters, and optimisation settings.
Figure 8: Impact of model size on forgetting dynamics. Plots illustrating the impact of varying numbers of hidden-layer parameters in a single-hidden-layer neural network on a regression task with 40 training datapoints (details in Appendix F). Left: Forgetfulness as a function of the number of updates $k$, showing how the learner’s propensity to forget evolves over update steps. Middle: Forgetfulness throughout training, highlighting changes in the learner’s propensity to forget over time. Right: Training loss curves, comparing learning dynamics across model sizes. We observe that forgetting dynamics are strongly influenced by model size: forgetting increases with model size until the learner’s parameter count approaches or exceeds the task’s effective size, at which point forgetting reduces again. These dynamics have a major impact on training efficiency, as shown in Figure 4.
Effect of model size.
Figure 8 examines how model size affects forgetting and learning efficiency on the regression task detailed in Appendix F. Each plot compares networks with different hidden-layer sizes, trained on 40 datapoints that span both under- and over-parameterised tasks. We find that forgetting generally increases with model size, peaking near the point where the number of parameters matches the task’s effective complexity, and then decreases again in the highly overparameterised task. This inverted trend mirrors the behaviour of training efficiency discussed in Figure 4.
Figure 9: Effect of the momentum parameter on forgetting dynamics. Plots illustrating the impact of varying momentum coefficients during training of a single-hidden-layer neural network on the sinusoid regression task. Left: Forgetfulness as a function of the number of updates $k$. Middle: Forgetfulness over the course of training. Right: Training loss curves for each momentum value. We observe that forgetting dynamics exhibit periodicity across updates, with the mean forgetting and the oscillation amplitude increasing as momentum increases. Optimal training efficiency occurs at a momentum of 0.9, corresponding to an optimal balance of adaptation and stability.
Effect of momentum.
In Figure 9, we vary the momentum coefficient in stochastic gradient descent while training on the same regression task. The propensity to forget exhibits periodic fluctuations whose amplitude increases with higher momentum values. Both the mean level of forgetting and the oscillatory behaviour grow with momentum, indicating a trade-off between stability and adaptivity. The most efficient training occurs around a momentum of 0.9, where moderate levels of forgetting coincide with stable and fast convergence.
Effect of batch size.
Figure 10 studies the influence of batch size on forgetting dynamics. Smaller minibatches lead to higher update variability, resulting in larger oscillations and a greater propensity to forget as training progresses. As batch size increases, forgetting decreases and training becomes more stable, plateauing at around one-quarter of the training dataset size. Beyond this, forgetting remains non-zero and is very similar across batch sizes greater than one-quarter of the training dataset.
Figure 10: Effect of batch size on forgetting dynamics. Lines and markers indicate different batch sizes used in training a single-hidden-layer neural network on the sinusoid regression task. Left: Forgetfulness as a function of the number of updates $k$. Middle: Forgetfulness over the course of training. Right: Training loss curves across batch sizes. Smaller batches exhibit significant fluctuations and high variability, resulting in unstable learning and reduced efficiency. Increasing batch size stabilises learning and reduces forgetting, until the effect plateaus at a batch size of 10, beyond which forgetting remains non-zero but is approximately the same regardless of batch size.
Figure 11: Forgetting dynamics across model architectures on a high-dimensional classification task. Different-coloured lines and marker styles denote different model types: logistic regression, MLP, CNN, and ResNet. Left: Forgetfulness over training updates. Right: Validation accuracy over training time. While CNNs and ResNets exhibit substantially higher levels of forgetting than logistic regression and MLPs (with ResNets forgetting the most), they achieve the best task performance. This implies that the most effective deep learning models do not necessarily minimise forgetting.
Effect of architecture.
Finally, Figure 11 examines how model architecture influences forgetting on a high-dimensional classification task (CIFAR-10). We compare logistic regression, a multilayer perceptron (MLP), a convolutional neural network (CNN) and a ResNet. CNNs and ResNets exhibit considerably higher forgetting than logistic regression and MLPs, yet they achieve better task performance. This indicates that the most effective deep learning models do not necessarily minimise forgetting; instead, they maintain a balance where moderate forgetting supports continued adaptation and improved generalisation.
Summary.
Together, these results (Figure 8, Figure 9, and Figure 10) show that forgetting is a pervasive property of deep learning systems, shaped by both the update dynamics and the learner’s hyperparameters. While high levels of forgetting can destabilise training, moderate levels appear beneficial: enabling adaptability while preserving some past information. The relationship between forgetting and learning efficiency is thus nuanced: efficient learning requires not the absence of forgetting but its regulation.
–appendices continue on next page–
Appendix F Experiment Details
Hardware.
We used NVIDIA GeForce RTX 2080 Ti GPUs to run our experiments.
Hyperparameters.
Tables 4 and 3 lists all of the hyperparameters that were used in the experiments.
Open-source code.
Upon acceptance, our code will be made available on GitHub.
Table 3: Hyperparameters for generative modelling and DQN reinforcement learning experiments.
Hyperparameter Description Values / Type
Forgetfulness settings
k Number of forgetfulness updates 40
num_particles Number of particles for Monte Carlo approximation 1000
Class Incremental-Learning
setting Dataset used Two Moons
noise Dataset noise 0.1
epochs Number of training epochs 30
num_tasks Number of tasks 2
batch_size Batch size 25
num_samples Number of training samples per task 100
num_val_samples Number of validation samples 100
lr Learning rate 0.1
optimiser Optimiser Adam
hidden_dim Hidden layer size 10
Reinforcement Learning (DQN)
setting Environment CartPole
batch_size Number of transitions sampled per gradient update 128
buffer_size Replay buffer size 10,000
start_e Initial exploration rate ($\epsilon$) 1.0
end_e Final exploration rate ($\epsilon$) 0.05
exploration_fraction Fraction of training during which $\epsilon$ decays 0.5
learning_starts Number of steps before training starts 10,000
train_frequency Number of steps between gradient updates 10
target_network_frequency Frequency of hard target network updates 500
tau Soft target update rate 1.0
gamma Discount factor for future rewards 0.99
lr Learning rate for optimiser 0.00025
optimiser Optimiser used for training Adam
num_eval_steps Number of steps per evaluation 1,000
total_timesteps Total number of environment interactions 200,000
num_parameters Number of model parameters 5
Table 4: Summary of hyperparameters and settings for all tasks. Most hyperparameters are fixed within each task, with task-specific variations highlighted where relevant.
Hyperparameter Description Values / Type
Regression
setting Dataset used Sinusoid
noise Observation noise 0.1
epochs Number of training epochs 30
batch_size Batch size 10
num_samples Number of training samples 40
num_val_samples Number of validation samples 100
lr Learning rate 0.1
optimiser Optimiser Adam
num_parameters Number of model parameters 5
Classification
setting Dataset used Two Moons
noise Observation noise 0.1
epochs Number of training epochs 30
batch_size Batch size 25
num_samples Number of training samples 100
num_val_samples Number of validation samples 100
lr Learning rate 0.1
optimiser Optimiser Adam
hidden_dim Hidden layer size 10
Generative Modelling
setting Dataset used Two Moons
batch_size Number of samples per training batch 2,500
epochs Number of training epochs 250
hidden_dim Number of hidden units in the model 64
lr Learning rate for optimiser 0.01
optimiser Optimiser used for training Adam
noise Noise added to dataset 0.05
num_integration_steps Number of integration steps in the model 100
num_samples Total number of training samples 10,000
num_val_samples Number of validation samples 1,000